OpenStudy (anonymous):

A truck begins a trip of 1675 miles. The truck averages 55 miles per hour, including stops.

OpenStudy (anonymous):

Create an equation to represent the distance (d) that the truck has left to travel with relation to time (t). d(t) =

OpenStudy (anonymous):

[Rate]=(55 mi/hr) (Distance Traveled)= (Rate)*(Time) [Distance Left]=1675miles-[Distance Traveled] Therefore, [Distance Left]= d(time)= 1675miles-(55 miles/hr)(time) or if you don't need units =1675-55t

OpenStudy (anonymous):

can you help me with another @enelius ?

OpenStudy (anonymous):

Sure.

OpenStudy (anonymous):

Alex is painting a house. When Alex starts work on Monday morning, there are 2452 square feet of surface area that remain to be painted. Alex can paint 64 square feet of surface area in an hour. Write an equation for the amount of surface area (A) that Alex has left to paint after t hours. A(t) =

OpenStudy (anonymous):

@enelius

OpenStudy (anonymous):

Alrighty, same type of deal as the first. Work Rate= (64 ft^2/hr) Work Completed = (Work Rate /*or power*/)*(Time Working) Work Left = 2452 ft^2 - (Work completed) Work Left = 2452 ft^2 - (64ft^2/hr)*(Time) A(t)=2452ft^2-(64ft^2/hr)*(t)